OpenStudy (anonymous):

Where have I made an error? I understand that I'm over complicating here, but i was practicing for the sake of practicing rather than solving the problem.

OpenStudy (anonymous):

x''+8x'+7x=2e^(-t) Complimentary: x''+8x'+7x=0 (D^2+8D+7)x=0 (D+1)(D+7)x=0 (D+7)x=u (D+1)u=0 du/dt+u=0 u=Ae^(-t) dx/dt+7x=Ae^(-t) dx/dt+dh/dt x=Ae^(-t) Be^ht dx/dt+Be^ht dh/dt x=Be^ht Ae^(-t) (d(Be^ht x))/dt=Be^ht Ae^(-t) 〖Be〗^ht x=∫▒〖BAe^((h-1)t) dt〗 e^ht x=A(1/(h-1))e^((h-1)t) x=A(1/(h-1))e^t 7= dh/dt h=7t x=A(1/(7t-1))e^t

OpenStudy (anonymous):

〖Be〗^ht =B e^ht